OpenStudy (mimi_x3):

A cylindrical metal container with base radius of r metres and height h metres holds 54πm^3. (a) Show that the surface area of the container is πr^2 +(108π/r) no lid.

OpenStudy (mimi_x3):

Easy ? Then help me then xD

OpenStudy (mimi_x3):

Gosh , i h8 proving

OpenStudy (saifoo.khan):

S.A. of a cylinder = \[2\pi r + 2 \pi r h\]

OpenStudy (mimi_x3):

I know that , im not stupid btw ,I need to show argh

OpenStudy (saifoo.khan):

Find the Radius first.

OpenStudy (saifoo.khan):

Volume = pi r h 54 = r * h

OpenStudy (saifoo.khan):

Still there is something missing!

OpenStudy (mimi_x3):

What's missing ?>

OpenStudy (saifoo.khan):

Radius or Height.

OpenStudy (mimi_x3):

LOL, nvm i already got help for another website xD . Thank you anyway (: